If i have 100 gallons of X mixed with 40 gallons of Y. What is the percentage of Y (in my 140 gallons)
Answer by stanbon(57347)

If i have 100 gallons of X mixed with 40 gallons of Y. What is the percentage of Y (in my 140 gallons)
-----------
x/100 = 40/140
x/100 = 4/14
x = 400/14 = 28.57%
=======================
Cheers,
Stan H.